two parallel dipoles of dipole moments p1 and p2 are placed at a x distance apart.what is d force on one dipole due to other?

Grade:

1 Answers

Sangarshanan Veera
9 years ago
p1..............................p2
pZ= q(2dx)     force acts due to charges +q and -q on any one dipole on the other....here i assume it as p2
Fnet = fq-f-q
Fnet = q(E+dE) -q(E-de)  where E = 2kp/x3
fnet = 2qdE  which is 2p2/2Dx dE
F = p2(dE/dX)
F = p2 *2kp1/x3 dX
F = p2*2kp1(3/Xpower4)
where k has the same meaning